OpenStudy (anonymous):

simplify (2/x^2+8x+15) + (1/x^2+11x+30)

OpenStudy (anonymous):

First step: \[\frac{2}{(x+5)(x+3)}+\frac{1}{(x+6)(x+5)}\]

OpenStudy (anonymous):

Second step: \[\frac{2(x+6)+(x+3)}{(x+5)(x+3)(x+6)}\]

OpenStudy (anonymous):

I'll let you take over

OpenStudy (anonymous):

can you cross out the x+3 and the x+6? and get 2/x+5 ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

You want to distribute the 2 to the numerator and rewrite the numerator

OpenStudy (anonymous):

so you would get 4x+18?

OpenStudy (anonymous):

\[\frac{2(x+6)+(x+3)}{(x+5)(x+3)(x+6)}=\frac{2x+12+x+3}{(x+6)(x+5)(x+3)}\] \[=\frac{3x+15}{(x+6)(x+5)(x+3)}\] Can you do the next step?

OpenStudy (anonymous):

so you dont distribute the 2 to x+3?

OpenStudy (anonymous):

No, there's a plus sign in between

OpenStudy (anonymous):

oh gotcha

OpenStudy (anonymous):

You only distribute for multiplication

OpenStudy (anonymous):

There's only a little bit left to do

OpenStudy (anonymous):

combine (x+5)(x+3)(x+6)?

OpenStudy (anonymous):

Nope, it has to do with the numerator first

OpenStudy (anonymous):

can you factor anything out of the numerator?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

3(x+5)

OpenStudy (anonymous):

Right. Now you can cancel the (x-5) in the numerator with the one in the denominator

OpenStudy (anonymous):

You're left with: \[\frac{3}{(x+6)(x+3)}\]

OpenStudy (anonymous):

its an x+5 though is that okay?

OpenStudy (anonymous):

never mind they both are

OpenStudy (anonymous):

That's as far as it can be simplified. If you want to FOIL the denominator, that's up to you.

OpenStudy (anonymous):

thanks :)